# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v2

@@ V Reduced dimensionality of quasisymmetric B
where the roots of the cubic polynomial, $B_{i}=\{B_{M},B_{m},B_{X}\}$, are functions of $\psi$, and $\mathcal{D}(\psi)$ is a proportionality factor. The exact solution of (52) subject to the periodicity constraint (11) can be expressed in terms of elliptic functions [30, 9].
@@
$B$ is then obtained from the cnoidal solution [30] of KdV,
@@
Here, $\text{cn}(u,m)$ is the Jacobi cosine elliptic function with modulus $m$. $B$ varies between $B_{m}$ and $B_{M}$. The period of $B$, $L(\psi)$, diverges logarithmically as $m\to 1$. One would expect this behavior near a separatrix, where periodicity along a field line is lost because the field line cannot cross the X-point. In particular, it is interesting to note that in the $m\to 1$ limit, $B$ takes the form of the infinite KdV soliton (reflectionless potential). The adiabatic invariant and integrals of the form (17) obtained from KdV potentials are manifestly time ($\alpha$) independent [38, 39, 78].
@@
The functions $B_{M}$ and $B_{m}$ denote the maximum and minimum values of $B$ on a flux surface, respectively, and $B_{X}$ is a third root of the polynomial. The extrema of $B$, given by $B_{M}$ and $B_{m}$ depend only on $\psi$, which implies that they can be straightened out using straight field line coordinates. The behavior of $B$ in Boozer coordinates is well known [17, 18, 27] near the magnetic axis and is consistent with (52). On the axis $B$ is a constant, $B_{0}$, and varies as $B\approx B_{0}(1+\overline{\eta}\sqrt{2\psi}\cos\vartheta)+2\psi(B_{20}+B_{22}\cos 2\vartheta)$ near the axis. Here, $\vartheta$ is the well-known Boozer angle [27]. The constant $\overline{\eta}$ dominates the maximum variation of $B$ on a flux surface, with the $O(2\psi)$ term being only a small second-order correction. Within the NAE framework, $B_{X}$ and $\mathcal{D}$ are constant throughout the volume, with $\mathcal{D}\propto((\iota-N/M))^{2}$. The NAE predictions are
@@ VI Numerical verification
Figure 3: Prediction of maximum toroidal volume that can be made quasisymmetric in precise QA. The boundary of the extended precise QA and two other flux surfaces are shown inside a Poincare plot of the islands in (a), and the roots are plotted in (b). The boundary of the original precise QA is shown in red. Note that the islands occur where the second and third roots touch approximately.
@@
Figure 19: $(\partial_{l}B)^{2}$ is fit to a polynomial of $B$ across all flux surfaces – a cubic in all cases except c), which is fit to a quartic as per Fig. 14. Shown is the evolution of the roots.
We shall now demonstrate numerically that a large class of quasisymmetric $B$ satisfies the integrated form of KdV as given in (52). As demonstrated in Fig. 1(a), $(dB/dl)^{2}$, when plotted as a function of $B$, has the form of a cubic polynomial in the precise QA equilibrium [4]; the scatter is due to imperfect QS. On the other hand, Fig. 1(b) shows that increasing the polynomial order beyond three does not lead to an observable decrease in error (defined as $1-r^{2}$, where $r^{2}$ is the coefficient of determination), as the fit error is dominated by QS error. Thus, Fig. 1 supports equation (52) as a model for a device with perfect QS. A similar analysis for other configurations, which tend to have more QS error, is shown in Appendix E. (Note that $s$ is the normalized toroidal flux in all figures.)
@@
With the global description of the exact quasisymmetric $B$, (55), at our disposal, we now investigate a key element of this theory: the physical meaning of the roots $B_{i}$. In Fig. 3(b), we show the typical behavior of the roots. Symmetry-breaking modes have been filtered out, and the roots are obtained by a cubic fit of the numerical data. Points represent the actual minimum and maximum values of $B$ on each flux surface. Surprisingly, the NAE description of $B_{M}$ and $B_{m}$ continues to hold even far from the axis. The third root $B_{X}$ depends quite sensitively on the $\iota$ profiles (Fig. 4). We show in Fig. 4 that for small but finite shear, $B_{X}$ tracks the $\iota$ profile quite closely but varies more for positive shear. This is in accordance with our free-surface approach to QS, which showed that $B_{X}$ must be a function of the shear parameter.
@@
Next, we demonstrate how we can estimate the maximum volume that a quasisymmetric field can occupy only from $B$ properties for any geometry. From the analytical solution for $B$ (55), we find that the periodicity of $B$ is maintained as long as $10\leq m<$ 1. As we saw earlier, the limit $(m\to 1,B_{m}\to B_{X})$ was the limit in which the connection length diverges. Using the simple NAE estimates for the roots $B_{m}$ and $B_{X}$, we can estimate where the two roots will intersect, thereby signaling the breakdown of periodicity. We have demonstrated in Fig. 3 how this estimate can predict the maximum volume of nested surfaces in precise QA extended using coils developed recently [79]. The appearance of the island chain in Fig. 3 is due to the failure of QS (which guarantees nested flux surfaces [22]) and not due to a resonance with the helicity of the QA [80].
Lastly, we show how the integrated form of KdV is approached through the stages of optimization, where the parameter space is progressively expanded. In Fig. 6, for the Landreman-Paul precise QA configuration, not only does the ‘noise’ decrease (as expected from improving QS), but the shape gets increasingly cubic. We show this both for a surface close to the axis and a surface close to the edge. Figures 7 and 8 further show how the KdV integrated cubic increases in accuracy (as determined by the coefficient of determination) further into the optimization process. In Figures 9, 10, 11, and 12, we show similar plots for the low mean rotational transform Buller configuration; note that the trend is towards quartic instead of cubic. As evident from these plots, the PP is established in the entire volume within the very first few stages of optimization and is maintained thereafter.
@@ VII Data driven verification
We have put forth several analytical arguments in favor of the connection between QS and the well-known integrable PDEs, KdV, and Gardner’s equation (for small rotational transforms). Our analysis so far has been based on NAEs, an approximate slab geometry model calculation and its extensions to more realistic geometries, and finally, soliton theory. We have also demonstrated numerically good agreement with several configurations with good QS. In this Section, we take a different approach based on data-driven techniques to determine the underlying system of equations describing quasisymmetric $B$ on a flux surface. Our primary tool is pySINDy [34, 35], which fits for governing differential equations and favors fewer terms to combat over-fitting.
We have primarily used the family of quasi-axisymmetric VMEC equilibria generated in [36]. For additional verification, we considered further configurations from the QUASR database [37]. Henceforth, we shall refer to them as the Buller and the Giuliani configurations. For the latter, configurations were selected with three representative mean rotational transforms. They were chosen to have the number of field periods, $N_{fp}=2$, an aspect ratio close to six (to mirror the Buller configurations), and the lowest possible quasi-axisymmetry error (as defined in [37]).
@@
Next, we use pySINDy to discover the underlying ODEs that govern $B$ on a flux surface. To that end, a sequentially thresholded least squares algorithm [81, 82] was used for the fitting, with a cost function that promotes sparsity in the data fit. For each magnetic configuration’s flux surface, the $B$ against $l$ profiles of multiple field lines spanning the surface were used as training data. The quantity $(\partial_{\ell}B)^{2}$ on a flux surface was fit against a linear combination of terms from a ‘feature library,’ chosen here to be a selection of polynomial terms, with the model parameters as the coefficients. Since the KdV and Gardner’s equation, together with the traveling waveform, lead to cubic and quartic polynomials, using the polynomial basis in pySINDy was a natural choice. A range of models was produced at varying sparsity parameters; the models were then tested on data from magnetic field lines not used in the training data. For sparsity parameters that were too large, it was found that there was a sharp drop in the success of the fit as too many terms were dropped; we chose the pySINDy model before this drop in performance. This is a version of hyperparameter scanning to find the Pareto-optimal fit (optimal tradeoff of model complexity versus model accuracy) to the data.
@@ VIII Discussions
In summary, we have demonstrated that periodic soliton solutions of the KdV equation lead to a broad class of excellent quasisymmetric $B$. We have done so by employing three different approaches. In the first approach, we reformulate the QS problem as a free-surface problem. The inner (closer to the axis) solution is assumed to be the second-order NAE solution, which has been demonstrated to capture essential features of even low-aspect ratio practical designs. The outer free-surface solution is then constructed by considering the length scale over which the magnetic shear varies. We can perturbatively solve the overdetermination problem by ensuring compatibility of the force balance and the QS condition. Our results show that shear is constrained to have an optimum value by the overdetermination problem and the magnetic field strength. The field strength must also satisfy the KdV equation, a well-known integrable PDE. We can also obtain Gardner’s equation for small values of the rotational transform.
@@
Firstly, the description of a class of precisely quasisymmetric $B$ can be reduced to understanding the behavior of the roots $B_{i}$ of a cubic polynomial. These roots correspond to extrema or saddle points of $B$ on a surface. In particular, $B$, which satisfies the periodic KdV equation, is the potential of the periodic Schrodinger equation where $B_{i}$’s are related to the endpoints of the band gap [30, 9]. Since the spectrum of KdV is time-independent, the roots do not depend on time. This confirms the well-known statement that the extrema or saddle points of $B$ depend only on the flux label in QS [11]. Moreover, the free-surface approach showed that $B_{X}$ is a function of the shear. Therefore, the invariance of $B_{X}$ on a flux surface effectively implies that the rotational transform $\iota(\psi)$ must be independent of the field line label (time). In other words, QS leads to an integrable field line Hamiltonian, which preserves nested flux surfaces [11]. Furthermore, since various aspects of the flux surface shaping can be deduced [12] from the Fourier coefficients of $B$, the hidden lower dimensionality is beneficial for understanding shaping. A similar hidden lower dimensionality is found in the description of the curvature of isodynamic magnetic field lines [28]. Once again, we emphasize that these hidden lower dimensionalities in the field strength or geometry are a direct consequence of the underlying soliton dynamics: the KdV equation in the case of QS and VFE in the case of isodynamic fields.
Secondly, the hierarchy of quasisymmetric invariants, as given by (17), can be understood in terms of the hierarchy of the conserved quantities of the KdV equation. Evolution under the KdV equation guarantees the field-line independence of these quantities. Thirdly, we can estimate the maximum toroidal volume QS allows solely from the properties of the roots and independent of any particular geometry. The overlap of the third root with the second signals the breakdown of periodicity of $B$, as demonstrated numerically. Finally, the connection length can be much larger than the major radius, typically assumed in the literature. A longer connection length may have important implications for linear and nonlinear stability in devices with QS.
@@
Our third approach is data-driven. Utilizing the vast amount of data on quasisymmetric equilibria generated in [36] and the QUASR database [37], and the state-of-the-art tool pySINDy, which uses sparse regression techniques to discover underlying models, we recover KdV and Gardner’s equation. Thus, the Painlevé property that demands that $(\partial_{\ell}B)^{2}$ can be at most a quartic in $B$ is verified accurately. This data-driven approach has the advantage that it can be applied to any configuration (including when quasi-symmetry is not relevant) to find low-dimensional ODE or PDE models underlying the data.
@@
Given the well-known robust stability of soliton solutions [83], the spectral stability of the cnoidal waves [84], and the robustness of the single-valuedness criterion, which led to the Painlevé property, perhaps explain why any numerical optimizer trying to find good QS using the 2-term form or the triple product form will most likely find only cubic or quartic terms in $(\partial_{\ell}B)^{2}$.
Acknowledgments

### v3

@@ V Reduced dimensionality of quasisymmetric B
where the roots of the cubic polynomial, $B_{i}=\{B_{M},B_{m},B_{X}\}$, are functions of $\psi$, and $\mathcal{D}(\psi)$ is a proportionality factor. The exact solution of (52) subject to the periodicity constraint (11) can be expressed in terms of elliptic functions (Kamchatnov, 2000; Novikov et al., 1984).
@@
$B$ is then obtained from the cnoidal solution (Kamchatnov, 2000) of KdV,
@@
Here, $\text{cn}(u,m)$ is the Jacobi cosine elliptic function with modulus $m$. $B$ varies between $B_{m}$ and $B_{M}$. The period of $B$, $L(\psi)$, diverges logarithmically as $m\to 1$. One would expect this behavior near a separatrix, where periodicity along a field line is lost because the field line cannot cross the X-point. In particular, it is interesting to note that in the $m\to 1$ limit, $B$ takes the form of the infinite KdV soliton (reflectionless potential). The adiabatic invariant and the integrals (17) obtained from KdV potentials are manifestly time ($\alpha$) independent (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990; Brezhnev, 2008).
@@
The functions $B_{M}$ and $B_{m}$ denote the maximum and minimum values of $B$ on a flux surface, respectively, and $B_{X}$ is a third root of the polynomial. The extrema of $B$, given by $B_{M}$ and $B_{m}$ depend only on $\psi$, which implies that they can be straightened out using straight field line coordinates. The behavior of $B$ in Boozer coordinates is well known (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2018) near the magnetic axis and is consistent with (52). On the axis $B$ is a constant, $B_{0}$, and varies as $B\approx B_{0}(1+\overline{\eta}\sqrt{2\psi}\cos\vartheta)+2\psi(B_{20}+B_{22}\cos 2\vartheta)$ near the axis. Here, $\vartheta$ is the well-known Boozer angle (Landreman and Sengupta, 2018). The constant $\overline{\eta}$ dominates the maximum variation of $B$ on a flux surface, with the $O(2\psi)$ term being only a small second-order correction. Within the NAE framework, $B_{X}$ and $\mathcal{D}$ are constant throughout the volume, with $\mathcal{D}\propto((\iota-N/M))^{2}$. The NAE predictions are
@@
We now look at several possibilities to understand the physical mechanisms underlying the divergence of the connection length. Broadly, these scenarios can be understood as the breakdown of the global straight field-line coordinates, the most important case being the breakdown of flux surfaces. Another possible scenario is the formation of magnetic puddles (Rodriguez and Plunk, 2023).
When the rotational transform is rational, X-points can form, near which the connection length diverges. X-points lead to the breakdown of the nestedness of flux surfaces and the formation of islands. If QS is exact, X-points can form only when $\iota$ resonates with the helicity of QS (Rodríguez and Bhattacharjee, 2021). However, even for minor QS errors, small resonant magnetic fields can lead to the formation of X-points and islands. For irrational $\iota$, sharp structures on the flux surfaces (”ridges”) can also arise (Bader et al., 2019; Punjabi and Boozer, 2020). These sharp structures can hinder magnetic field lines from crossing them (Boozer, 2015). Finally, topological obstructions such as the “Reeb components” (Duignan et al., 2024) can also lead to the non-existence of global straight field-line coordinates. However, the Reeb case requires the toroidal magnetic field to vanish at some point if MHS force balance is imposed (Duignan et al., 2024). Similarly, magnetic field strength topology breakdown can be avoided if QS does not change drastically or by numerically filtering out any non-QS part of $B$.
In Section VI, we indeed observe the formation of an X-point as we approach the limiting flux surface. In practical terms, this means that QS stellarators may naturally form sharp edges as the volume of quasisymmetry is pushed to its limit, e.g., for compact devices. Sharp edges are essential for non-resonant divertors (Boozer, 2015; Punjabi and Boozer, 2020). This could have significant beneficial consequences for stellarator optimization, since these divertors would essentially be obtained for free by optimizing for quasisymmetry in compact devices.
@@ VI Numerical verification
Figure 19: $(\partial_{l}B)^{2}$ is fit to a polynomial of $B$ across all flux surfaces – a cubic in all cases except a), which is fit to a quartic as per Fig. 14. Shown is the evolution of the roots.
We shall now demonstrate numerically that a large class of quasisymmetric $B$ satisfies the integrated form of KdV as given in (52). As demonstrated in Fig. 1(a), $(dB/dl)^{2}$, when plotted as a function of $B$, has the form of a cubic polynomial in the precise QA equilibrium Landreman and Paul (2022); the scatter is due to imperfect QS. On the other hand, Fig. 1(b) shows that increasing the polynomial order beyond three does not lead to an observable decrease in error (defined as $1-r^{2}$, where $r^{2}$ is the coefficient of determination), as the fit error is dominated by QS error. Thus, Fig. 1 supports equation (52) as a model for a device with perfect QS. A similar analysis for other configurations, which tend to have more QS error, is shown in Appendix E. (Note that $s$ is the normalized toroidal flux in all figures.)
@@
With the global description of the exact quasisymmetric $B$, (55), at our disposal, we now investigate a key element of this theory: the physical meaning of the roots $B_{i}$. In Fig. 3(b), we show the typical behavior of the roots. Symmetry-breaking modes have been filtered out, and the roots are obtained by a cubic fit of the numerical data. Points represent the actual minimum and maximum values of $B$ on each flux surface. Surprisingly, the NAE description of $B_{M}$ and $B_{m}$ continues to hold even far from the axis. As shown in Fig. 3(a), the divergence in the connection length as the two roots meet is indeed due to the formation of an X-point that leads to a chain of islands.
The third root $B_{X}$ depends quite sensitively on the $\iota$ profiles (Fig. 4). We show in Fig. 4 that for small but finite shear, $B_{X}$ tracks the $\iota$ profile quite closely but varies more for positive shear. This is in accordance with our free-surface approach to QS, which showed that $B_{X}$ must be a function of the shear parameter.
@@
Next, we demonstrate that the maximum volume that a quasisymmetric field can occupy can be estimated by the crossings of the roots of $\partial B/\partial l$ (52). From the analytical solution for $B$ (55), we find that the periodicity of $B$ is maintained as long as $10\leq m<$ 1. As we saw earlier, the limit $(m\to 1,B_{m}\to B_{X})$ was the limit in which the connection length diverges. Using the simple NAE estimates for the roots $B_{m}$ and $B_{X}$, we can estimate where the two roots will intersect, thereby signaling the breakdown of periodicity. We have shown in Fig. 3 how this estimate can predict the maximum volume of nested surfaces in precise QA extended using coils developed recently Wechsung et al. (2022). We note that the appearance of the resonant 4/10 island chain in Fig. 3 is not due to a resonance with the helicity of the QA Rodríguez and Bhattacharjee (2021). We further note that since the non-QS modes were filtered out, the breakdown of the topology of the field strength does not play any role in the divergence of the connection length.
Lastly, we show how the integrated form of KdV is approached through the stages of optimization, where the parameter space is progressively expanded. In Fig. 6, for the Landreman-Paul precise QA configuration, not only does the ‘noise’ decrease (as expected from improving QS), but the shape gets increasingly cubic. We show this both for a surface close to the axis and a surface close to the edge. Figures 7 and 8 further show how the KdV integrated cubic increases in accuracy (as determined by the coefficient of determination) further into the optimization process. In Figures 9, 10, 11, and 12, we show similar plots for the low mean rotational transform configurations introduced in Buller et al. (2024). Note that the trend is towards quartic instead of cubic. As evident from these plots, the PP is established in the entire volume within the very first few stages of optimization and is maintained thereafter.
@@ VII Data driven verification
We have put forth several analytical arguments in favor of the connection between QS and the well-known integrable PDEs, KdV, and Gardner’s equation (for small rotational transforms). Our analysis so far has been based on NAEs, an approximate slab geometry model calculation and its extensions to more realistic geometries, and finally, soliton theory. We have also demonstrated numerically good agreement with several configurations with good QS. In this Section, we take a different approach based on data-driven techniques to determine the underlying system of equations describing quasisymmetric $B$ on a flux surface. Our primary tool is pySINDy de Silva et al. (2020); Kaptanoglu et al. (2022), which fits for governing differential equations and favors fewer terms to combat over-fitting.
We have primarily used the family of quasi-axisymmetric VMEC equilibria generated in Buller et al. (2024). For additional verification, we considered further configurations from the QUASR database Giuliani (2023). Henceforth, we shall refer to them as the Buller and the Giuliani configurations. For the latter, configurations were selected with three representative mean rotational transforms. They were chosen to have the number of field periods, $N_{fp}=2$, an aspect ratio close to six (to mirror the Buller configurations), and the lowest possible quasi-axisymmetry error (as defined in Giuliani (2023)).
@@
Next, we use pySINDy to discover the underlying ODEs that govern $B$ on a flux surface. To that end, a sequentially thresholded least squares algorithm Brunton et al. (2016); Zhang and Schaeffer (2019) was used for the fitting, with a cost function that promotes sparsity in the data fit. For each magnetic configuration’s flux surface, the $B$ against $l$ profiles of multiple field lines spanning the surface were used as training data. The quantity $(\partial_{\ell}B)^{2}$ on a flux surface was fit against a linear combination of terms from a ‘feature library,’ chosen here to be a selection of polynomial terms, with the model parameters as the coefficients. Since the KdV and Gardner’s equation, together with the traveling waveform, lead to cubic and quartic polynomials, using the polynomial basis in pySINDy was a natural choice. A range of models was produced at varying sparsity parameters; the models were then tested on data from magnetic field lines not used in the training data. For sparsity parameters that were too large, it was found that there was a sharp drop in the success of the fit as too many terms were dropped; we chose the pySINDy model before this drop in performance. This is a version of hyperparameter scanning to find the Pareto-optimal fit (optimal tradeoff of model complexity versus model accuracy) to the data.
@@ VIII Discussions
In summary, we have demonstrated that periodic soliton solutions of the KdV equation lead to a broad class of excellent quasisymmetric $B$. We have done so by employing three different approaches. In the first approach, we reformulate the QS problem as a free-surface problem. The inner (closer to the axis) solution is assumed to be the second-order NAE solution, which has been demonstrated to capture essential features of even low-aspect-ratio practical designs. The outer free-surface solution is then constructed by considering the length scale over which the magnetic shear varies. We can perturbatively solve the overdetermination problem by ensuring compatibility of the force balance and the QS condition. Our results show that shear is constrained to have an optimum value by the overdetermination problem and the magnetic field strength. The field strength must also satisfy the KdV equation, a well-known integrable PDE. We can also obtain Gardner’s equation for small values of the rotational transform.
@@
Firstly, the description of a class of precisely quasisymmetric $B$ can be reduced to understanding the behavior of the roots $B_{i}$ of a cubic polynomial. These roots correspond to extrema or saddle points of $B$ on a surface. In particular, $B$, which satisfies the periodic KdV equation, is the potential of the periodic Schrodinger equation where $B_{i}$’s are related to the endpoints of the band gap (Kamchatnov, 2000; Novikov et al., 1984). Since the spectrum of KdV is time-independent, the roots do not depend on time. This confirms the well-known statement that the extrema or saddle points of $B$ depend only on the flux label in QS Helander (2014). Moreover, the free-surface approach showed that $B_{X}$ is a function of the shear. Therefore, the invariance of $B_{X}$ on a flux surface effectively implies that the rotational transform $\iota(\psi)$ must be independent of the field line label (time). In other words, QS leads to an integrable field line Hamiltonian, which preserves nested flux surfaces (Helander, 2014). Furthermore, since various aspects of the flux surface shaping can be deduced (Rodriguez, 2022) from the Fourier coefficients of $B$, the hidden lower dimensionality is beneficial for understanding shaping. A similar hidden lower dimensionality is found in the description of the curvature of isodynamic magnetic field lines (Schief, 2003). We repeat for mphasis that these hidden lower dimensionalities in the field strength or geometry are a direct consequence of the underlying soliton dynamics: the KdV equation in the case of QS and vortex filament equation in the case of isodynamic fields.
Secondly, the hierarchy of quasisymmetric invariants, as given by (17), can be understood in terms of the hierarchy of the conserved quantities of the KdV equation. Evolution under the KdV equation guarantees the field-line independence of these quantities.
Thirdly, we can estimate the maximum toroidal volume QS allows solely from the properties of the roots and independent of any particular geometry. The overlap of the third root with the second signals the breakdown of periodicity of $B$, as demonstrated numerically. At this radius, a separatrix could form, which could potentially be used as the basis for a non-resonant divertor (Boozer, 2015; Punjabi and Boozer, 2020).
Finally, the connection length can be much larger than the major radius, typically assumed in the literature. A longer connection length may have important implications for linear and nonlinear stability in devices with QS.
@@
Our third approach is data-driven. Utilizing the vast amount of data on quasisymmetric equilibria generated in Buller et al. (2024) and the QUASR database Giuliani (2023), and the state-of-the-art tool pySINDy, which uses sparse regression techniques to discover underlying models, we recover KdV and Gardner’s equation. Thus, the Painlevé property that demands that $(\partial_{\ell}B)^{2}$ can be at most a quartic in $B$ is verified accurately. This data-driven approach has the advantage that it can be applied to any configuration (including when quasi-symmetry is not relevant) to find low-dimensional ODE or PDE models underlying the data.
@@
Given the well-known robust stability of soliton solutions Tao (2009), the spectral stability of the cnoidal waves (Bottman and Deconinck, 2009), and the robustness of the single-valuedness criterion, which led to the Painlevé property, perhaps explain why any numerical optimizer trying to find good QS using the 2-term form or the triple product form will most likely find only cubic or quartic terms in $(\partial_{\ell}B)^{2}$.
Acknowledgements.